# Crucial-Diff: A Unified Diffusion Model for Crucial Image and Annotation Synthesis in Data-scarce Scenarios

### v1

@@ Abstract
The scarcity of data in various scenarios, such as medical, industry and autonomous driving, leads to model overfitting and dataset imbalance, thus hindering effective detection and segmentation performance. Existing studies employ the generative models to synthesize more training samples to mitigate data scarcity. However, these synthetic samples are repetitive or simplistic and fail to provide “crucial information” that targets the downstream model’s weaknesses. Additionally, these methods typically require separate training for different objects, leading to computational inefficiencies. To address these issues, we propose Crucial-Diff, a domain-agnostic framework designed to synthesize crucial samples. Our method integrates two key modules. The Scene Agnostic Feature Extractor (SAFE) utilizes a unified feature extractor to capture target information. The Weakness Aware Sample Miner (WASM) generates hard-to-detect samples using feedback from the detection results of downstream model, which is then fused with the output of SAFE module. Together, our Crucial-Diff framework generates diverse, high-quality training data, achieving a pixel-level AP of 83.63% and an F1-MAX of 78.12% on MVTec. On polyp dataset, Crucial-Diff reaches an mIoU of 81.64% and an mDice of 87.69%. Code will be released after acceptance.
Index Terms:
Scarce dataset generation, crucial sample generation, downstream model feedback, diffusion model.
@@ I Introduction
The scarcity of training samples poses significant challenges in developing reliable deep-learning models, such as the scenarios of the industrial quality inspection with rare defect images [1, 2, 3, 4, 5] and the medical image diagnosis with rare precise annotations [6, 7, 8, 9, 10]. Therefore, this study focuses on target segmentation tasks in these data-scarce situations, aiming to investigate methods for generating effective samples to alleviate data insufficiency. To address this challenge, the Images and Annotations Synthesis (IAS) task in data-scarce scenarios has emerged as a pivotal paradigm [11, 12]. It synthesizes diverse image samples with accurate segmentation annotations to expand limited data, enabling large-scale synthetic data to train and improve downstream models.
Early IAS methods in data-scarce scenarios relied on traditional data augmentation techniques [13], expanding the dataset through transformations like rotation and cropping, etc. In contrast, the latest generative approaches, like Generative Adversarial Networks (GANs) [14, 15] and diffusion models [11, 16], are employed to synthesize new images and annotations by simulating the data distribution of real samples. However, these synthesized samples, which we define as easy synthetic images, often exhibit distributions that closely overlap with real samples. As shown in Fig. 1(a), easy synthetic samples (yellow points) cluster tightly around real samples (pink points), demonstrating their limited diversity. This limitation is further highlighted in Fig. 1(b) of the industrial inspection scenario, where easy synthetic images appear visually identical to real anomalies, failing to introduce meaningful variations or challenges for downstream models.
As shown in the yellow curve of Fig. 1(c), only using the easy synthetic samples (the yellow points in Fig. 1(a)) initially boosts downstream task performance. However, this improvement quickly plateaus and may even decline as the data volume increases. This stagnation occurs because easy synthetic images lack diversity and cluster around simple, known patterns, leading downstream models to overfit on familiar data and generalize poorly to unseen scenarios. In contrast, crucial synthetic images, as seen in the purple points of Fig. 1(a), are samples that lie at hard detectable regions of downstream models, exposing their weaknesses and effectively improving their decision-making accuracy.
Additionally, significant domain gaps between different targets often necessitate training separate generative models for each object [17]. This approach consumes substantial training resources as the target number increases, making the acquisition of synthetic samples highly inefficient.
@@
To address these requirements, we propose Crucial-Diff, a domain-agnostic framework that focuses on crucial synthetic sample generation. By employing a diffusion model conditioned on the reference image of the target, Crucial-Diff synthesizes realistic targets on a background image without any target, positioned according to the specified box location, thereby eliminating the fine-tuning needs for different targets or scenarios. Two additional modules are proposed to address crucial synthetic sample generation and multi-scenario universality, respectively, to ensure that the generated samples are effectively bridge data scarcity.
@@ II-B Rare Image and Annotation Synthesis
Traditional approaches typically employ GANs to generate image-annotation pairs [32, 33], while contemporary diffusion models demonstrate superior capabilities for image-annotation synthesis [34, 35, 36]. DSMA [37] further enables adversarial attacks using synthetic dataset. However, these methods fail in extreme few-shot scenarios, particularly in industrial and medical applications. Therefore, we focus on developing generation models for these two fields.
@@ III Method
Crucial-Diff tackles target synthesis challenges in data-scarce scenarios by generating high-quality samples that expose downstream model weaknesses and ensure cross-domain applicability. As shown in Fig. 2(a), our Crucial-Diff synthesizes crucial images by inputting a target reference image $x^{r}$ that encodes target-specific features, a background image $x^{b}$ that provides contextual information and a bounding box mask $m$ that specifies the location of target into model. Specifically, the Scene Agnostic Feature Extractor module is first used to extract textual embeddings of the target characteristic from the reference image (Sec. III-B). Next, the Weakness Aware Sample Miner (WASM) extracts textual embeddings of subtle target characteristics that can mislead downstream model detection errors from the reference image, leveraging its ability to capture deceptive features acquired during the training process (Sec. III-C). The textual features from WASM module are used to compute attention outputs, which are then fused with those from SAFE module, dynamically guiding synthesis process to obtain crucial images with target. Finally, accurate pixel-level annotations of targets are produced by extracting and binarizing attention maps between crucial features and target features derived from proposed two modules (Sec. III-D).
@@ III-A Basic Pipeline
$\mathcal{L}=\mathbb{E}_{z_{t}^{r},p,t,\epsilon}\big{[}\big{|}\big{|}\epsilon-\epsilon_{\theta}(z_{t}^{r},p,t)\big{|}\big{|}^{2}\big{]}.$ (2)
@@ III-B Scene Agnostic Feature Extractor Module
$\mathcal{L}^{u}=\mathbb{E}_{z_{t}^{r},p^{u},t,\epsilon}\big{[}\big{|}\big{|}\epsilon-\epsilon_{\theta}(z_{t}^{r},p^{u},t)\big{|}\big{|}^{2}\big{]}+||p^{u}||_{1},$ (7)
@@ III-C Weakness Aware Sample Miner Module
In stage 2 of Fig. 2(b), the WASM module takes a real scarce image $x^{r}$ as input and extracts multi-layer visual features using the CLIP encoder $\Psi(\cdot)$. Unlike SAFE, which simply maps images to descriptions about targets, WASM module employs a target transfer network $\phi^{c}(\cdot)$ that shares the same architecture as the Target Extractor but serves a distinct purpose. Through subsequent fine-tuning, $\phi^{c}(\cdot)$ specializes in extracting discriminative textual embeddings from target images, which explicitly encode the subtle but critical features that downstream models most frequently misclassify.
@@
$Attn\leftarrow\lambda\cdot Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{c},\mathbf{V}_{t}^{c})+(1-\lambda)Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{u},\mathbf{V}_{t}^{u}),$ (8)
@@
$\mathcal{L}^{c}$ $=||\frac{\sum_{(i,j)\in m}f_{i,j}}{|m|+\delta}-\frac{\sum_{(i,j)\in(1-m)}f_{i,j}}{|1-m|+\delta}||^{2}+||p^{c}||_{1},$ (9)
where $|m|$ is the number of pixels in the bounding box and $|1-m|$ is the number of pixels outside it, and $\delta$ is a small constant to avoid division by zero. Notably, we first train SAFE module, using its parameters to initialize WASM module for stable training. The WASM module is then fine-tuned to synthesise crucial targets. This two-stage method is essential because joint optimization creates competing objectives. SAFE’s real target generation conflicts with WASM’s undetectable generation, ultimately reducing the effectiveness of the synthetic crucial targets.
@@ III-D Target Annotation Generation
Figure 3: Visualization of iterative refinement of the cross-attention layer. The first row displays cross-attention maps for iterative counts $I=0,1,3,5,7$, while the second row shows the corresponding pixel-level annotations $\hat{m}$.
The direct application of cross-attention maps frequently results in insufficient fine-grained spatial details, which can compromise the accuracy of target region localization when these maps are binarized for detailed annotations, particularly in complex scenes characterized by cluttered backgrounds. By harnessing the rich spatial information encoded in self-attention maps, we can significantly enhance the spatial precision of the attention maps, thereby facilitating a more accurate delineation of target regions. Similar to [49], we extract self-attention maps $\mathcal{A}^{sa}$ and cross-attention maps $\mathcal{A}^{ca}$ from multiple layers of the U-Net decoder and utilize the self-attention maps to iteratively refine the cross-attention maps. In each iteration $i$, the refined cross-attention map is continuously updated as follows:
@@
where $I$ is the total number of iterations. Notably, $\mathcal{A}^{ca}_{0}$ is the average of cross-attention maps activated by $p^{*}$ at last timestep. The normalization operation is applied after each iteration to ensure the stability of the refined map. As shown in Fig. 3, the refinement process progressively enhances the spatial accuracy of the attention map. However, excessive iteration may lead to overfitting of local textures while ignoring global structural patterns. We therefore select $I=5$ as the optimal balance. Finally, by applying a threshold to the refined cross-attention map, we effectively separate the target region from the background, resulting in pixel-level annotations $\hat{m}$ aligned with synthesized targets.
@@ IV-A Implementation Details
Figure 4: Visualization comparison of synthetic images generated by Crucial-Diff and other methods. (a) is MVTec result and (b) is polyp dataset result with the annotations of synthetic targets at bottom right of each image. Our generated targets blend seamlessly into the background, making distinguishing them challenging.
@@ IV-B Datasets and Metrics
Industrial inspection. We evaluate Crucial-Diff on the MVTec dataset [50], which contains 3,629 normal training images and a test set of 467 normal and 1,258 anomaly images with mask labels. Following [11], one-third of the anomaly images are used for training, while the rest is used to test downstream tasks. Generation performance is evaluated with Intra-cluster pairwise LPIPS distance (IC-LPIPS) [51], Inception Score (IS) [52] and Kernel Inception Distance (KID) [53]. Detection performance is evaluated with Average Precision (AP) and maximum F1 score (F1-MAX). Computational efficiency is measured by metric Time, which means total training hours for all targets.
@@
Figure 5: T-SNE visualization of different targets. The pink triangle represents a real image with target, the green circle represents a real image without target, the yellow pentagram represents a easy synthetic image with target, and the purple diamond represents a crucial synthetic image with target.
@@ IV-C Qualitative Comparison
As shown in Fig. 4(a), in MVTec dataset, previous methods, such as AnomalyDiffusion [11], tend to generate targets that are visually distinct and easily recognizable, limiting their effectiveness in improvement of downstream models. In contrast, Crucial-Diff excels at generating subtle anomalies, such as faint cuts and barely visible breaks. These subtle targets are easily overlooked by downstream tasks as training data can effectively improve detection results.
As shown in Fig. 4(b), in polyp the dataset, Crucial-Diff synthesizes polyps characterized by blurry shapes and low contrast, allowing them to blend naturally into the surrounding tissue. This enhances visual realism and improves model applicability in clinical settings. In contrast, previous methods often generate polyps with sharp edges and high contrast, making them easier to detect but less realistic.

### v2

@@ Abstract
The scarcity of data in various scenarios, such as medical, industry and autonomous driving, leads to model overfitting and dataset imbalance, thus hindering effective detection and segmentation performance. Existing studies employ the generative models to synthesize more training samples to mitigate data scarcity. However, these synthetic samples are repetitive or simplistic and fail to provide “crucial information” that targets the downstream model’s weaknesses. Additionally, these methods typically require separate training for different objects, leading to computational inefficiencies. To address these issues, we propose Crucial-Diff, a domain-agnostic framework designed to synthesize crucial samples. Our method integrates two key modules. The Scene Agnostic Feature Extractor (SAFE) utilizes a unified feature extractor to capture target information. The Weakness Aware Sample Miner (WASM) generates hard-to-detect samples using feedback from the detection results of downstream model, which is then fused with the output of SAFE module. Together, our Crucial-Diff framework generates diverse, high-quality training data, achieving a pixel-level AP of 83.63% and an F1-MAX of 78.12% on MVTec. On polyp dataset, Crucial-Diff reaches an mIoU of 81.64% and an mDice of 87.69%. Code is publicly available at https://github.com/JJessicaYao/Crucial-diff.
@@ I Introduction
The scarcity of training samples poses significant challenges in developing reliable deep-learning models, such as the scenarios of the industrial quality inspection with rare defect images [1, 2, 3, 4, 5, 6, 7] and the medical image diagnosis with rare precise annotations [8, 9, 10, 11, 12, 13, 14]. Therefore, this study focuses on target segmentation tasks in these data-scarce situations, aiming to investigate methods for generating effective samples to alleviate data insufficiency. To address this challenge, the Images and Annotations Synthesis (IAS) task in data-scarce scenarios has emerged as a pivotal paradigm [15, 16, 17, 18, 19]. It synthesizes diverse image samples with accurate segmentation annotations to expand limited data, enabling large-scale synthetic data to train and improve downstream models.
Early IAS methods in data-scarce scenarios relied on traditional data augmentation techniques [20], expanding the dataset through transformations like rotation and cropping, etc. In contrast, modern generative approaches, including Generative Adversarial Networks (GANs) [21, 22] and diffusion models [15, 17], aim to replicate training data distributions by minimizing statistical divergence between synthetic and real samples. As shown in Fig. 1(a), samples generated by these methods, which we refer to as “easy synthetic samples” (yellow points), cluster densely around real samples (pink points). This limitation is further highlighted in Fig. 1(b) of the industrial inspection scenario, where easy synthetic images appear visually identical to real anomalies, failing to introduce meaningful variations or challenges for downstream models.
As shown in the yellow curve of Fig. 1(c), only using the easy synthetic samples (the yellow points in Fig. 1(a)) initially boosts downstream task performance. However, this improvement quickly plateaus and may even decline as the data volume increases. This stagnation stems from the tendency of easy synthetic samples to cluster around simple, known patterns. While these samples quantitatively expand the dataset, they lack qualitative novelty or innovative features, causing downstream models to overfit on familiar patterns and exhibit poor generalization in unseen scenarios. In contrast, “crucial synthetic images” specifically designed to target uncertain and hard detectable feature of the downstream models, as shown by the purple points in Fig. 1(a). These samples expose the weaknesses of downstream models, significantly improving their accuracy by forcing the models to learn from these challenging samples.
Additionally, existing approaches such as DFMGAN [23] often train separate generative models for each target due to significant domain gaps between different targets. This per target training paradigm consumes substantial computational resources, resulting in exponentially increasing costs as the number of required targets grows. Such methods severely constrain the scalability of synthetic data, ultimately limiting their widespread adoption in real world applications.
@@
To address these requirements, we propose Crucial-Diff, a domain-agnostic framework that focuses on crucial synthetic sample generation. By employing a diffusion model conditioned on the reference image of the target, Crucial-Diff synthesizes realistic targets on a background image without any target, positioned according to the specified box location, thereby eliminating the fine-tuning needs for different targets or scenarios. In contrast to previous methods dependent on precise masks as location constraints for synthesis, which typically require additional mask generators, our method randomly generates bounding box masks without needing extra generators, significantly reducing computational overhead. Furthermore, two additional modules are proposed to address crucial synthetic sample generation and multi-scenario universality, respectively, to ensure that the generated samples are effectively bridge data scarcity.
@@ II-B Rare Image and Annotation Synthesis
Traditional approaches typically employ GANs to generate image-annotation pairs [37, 38], while contemporary diffusion models demonstrate superior capabilities for image-annotation synthesis [39, 40, 41, 42]. DSMA [43] further enables adversarial attacks using synthetic dataset. However, these methods fail in extreme few-shot scenarios, particularly in industrial and medical applications. Therefore, we focus on developing generation models for these two fields.
@@ III Method
Crucial-Diff tackles target synthesis challenges in data-scarce scenarios by generating high-quality samples that expose downstream model weaknesses and ensure cross-domain applicability. As shown in Fig. 2(a), our Crucial-Diff synthesizes crucial images by inputting a target reference image $x^{r}$ that encodes target-specific features, a background image $x^{b}$ that provides contextual information and a bounding box mask $m$ that specifies the location of target into model. The bounding box mask is generated by randomly sampling coordinates within the background image space, with both its position and size determined through a randomization process. Crucial-Diff begins with the Scene Agnostic Feature Extractor (SAFE) module, which extracts textual embeddings of the target characteristic from the reference image (Sec. III-B). Next, the Weakness Aware Sample Miner (WASM) extracts textual embeddings of subtle target characteristics that can mislead downstream model detection errors from the reference image, leveraging its ability to capture deceptive features acquired during the training process (Sec. III-C). The textual features from WASM module are used to compute attention outputs, which are then fused with those from SAFE module, dynamically guiding synthesis process to obtain crucial images with target. Finally, accurate pixel-level annotations of targets are produced by extracting and binarizing attention maps between crucial features and target features derived from proposed two modules (Sec. III-D).
@@ III-A Basic Pipeline
[…]
@@ III-B Scene Agnostic Feature Extractor Module
[…]
@@ III-C Weakness Aware Sample Miner Module
In stage 2 of Fig. 2(b), the WASM module takes a real scarce image $x^{r}$ as input and extracts multi-layer visual features using the CLIP encoder $\Psi(\cdot)$. Unlike SAFE, which simply maps images to descriptions about targets, WASM module employs a target transfer network $\phi^{c}(\cdot)$ that shares the same architecture as the target extractor network but serves a distinct purpose. Through subsequent fine-tuning, $\phi^{c}(\cdot)$ specializes in extracting discriminative textual embeddings from target images, which explicitly encode the subtle but critical features that downstream models most frequently misclassify.
@@
$Attn\leftarrow\lambda\cdot Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{c},\mathbf{V}_{t}^{c})+(1-\lambda)\cdot Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{u},\mathbf{V}_{t}^{u}),$ (8)
@@
[…]
where $|m|$ is the number of pixels in the bounding box and |1m||1-m| is the number of pixels outside it, and $\delta$ is a small constant to avoid division by zero. Notably, we first train SAFE module, using its parameters to initialize WASM module for stable training. The WASM module is then fine-tuned to synthesise crucial targets. This two-stage method is essential because joint optimization creates competing objectives. SAFE’s real target generation conflicts with WASM’s undetectable generation, ultimately reducing the effectiveness of the synthetic crucial targets.
@@ III-D Target Annotation Generation
Figure 3: Visualization of iterative refinement of the cross-attention layer. The first row displays cross-attention maps when the iteration is 0, 1, 3, 5 and 7, while the second row shows the corresponding pixel-level annotations $\hat{m}$.
During inference, the direct application of cross-attention maps frequently results in insufficient fine-grained spatial details, which can compromise the accuracy of target region localization when these maps are binarized for detailed annotations, particularly in complex scenes characterized by cluttered backgrounds. By harnessing the rich spatial information encoded in self-attention maps, we can significantly enhance the spatial precision of the attention maps, thereby facilitating a more accurate delineation of target regions. Similar to [55], we extract self-attention maps $\mathcal{A}^{sa}$ and cross-attention maps $\mathcal{A}^{ca}$ from multiple layers of the U-Net decoder and utilize the self-attention maps to iteratively refine the cross-attention maps. In each iteration $i$, the refined cross-attention map is continuously updated as follows:
@@
where $I$ is the total number of iterations. Notably, $\mathcal{A}^{ca}_{0}$ is the average of cross-attention maps activated by $p^{*}$ at last timestep. The normalization operation is applied after each iteration to ensure the stability of the refined map. As shown in Fig. 3, the refinement process progressively enhances the spatial accuracy of the attention map. However, excessive iteration may lead to overfitting of local textures while ignoring global structural patterns. We therefore select $I=5$ as the optimal balance. Finally, by applying a threshold to the refined cross-attention map, we effectively separate the target region from the background, resulting in pixel-level annotations $\hat{m}$ aligned with synthesized targets. Overall, the detailed synthesis process of images and annotations is depicted in Algo. 1.
@@ IV-A Implementation Details
Algorithm 1 The synthesis process of proposed Crucial-Diff.
Input: target image $x^{r}$, background image $x^{b}$, coarse mask $m$, total inference timestep $T$
Parameter: conditional noise predictor $\epsilon_{\theta}(\cdot)$, VAE image encoder $\varepsilon(\cdot)$, CLIP text encoder $\tau(\cdot)$, CLIP visual encoder $\Psi(\cdot)$, target extractor network $\phi^{u}(\cdot)$, target transfer network $\phi^{c}(\cdot)$ and hyperparameter $\lambda$
Output: the synthetic image $x^{s}$, corresponding annotation $\hat{m}$
[…]
2: Select $\Gamma$ textual embeddings from $\phi^{u}(\Psi(x^{r}))$.
[…]
5:for $t=T$ to 11 do
6:  for each cross attention layer do
7:   Compute $\mathbf{K}_{t}^{u}$, $\mathbf{V}_{t}^{u}$ using SAFE module.
8:   $\mathbf{K}_{t}^{u}=(\mathcal{W}_{t}^{u})^{K}\cdot p^{u}$;  $\mathbf{V}_{t}^{u}=(\mathcal{W}_{t}^{u})^{V}\cdot p^{u}$
9:   $Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{u},\mathbf{V}_{t}^{u})=\mathrm{softmax}(\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}^{u}}^{\top}{}}{\sqrt{d}})\cdot\mathbf{V}^{u}_{t}$
10:   Compute $\mathbf{K}_{t}^{c}$, $\mathbf{V}_{t}^{c}$ using WASM module.
11:   $\mathbf{K}_{t}^{c}=(\mathcal{W}_{t}^{c})^{K}\cdot p^{c}$;  $\mathbf{V}_{t}^{c}=(\mathcal{W}_{t}^{c})^{V}\cdot p^{c}$
12:   $Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{c},\mathbf{V}_{t}^{c})=\mathrm{softmax}(\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}^{c}}^{\top}{}}{\sqrt{d}})\cdot\mathbf{V}^{c}_{t}$
13:   Update $Attn\leftarrow\lambda\cdot Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{c},\mathbf{V}_{t}^{c})+(1-\lambda)\cdot Attn(\mathbf{Q}_{t},\mathbf{K}_{t}^{u},\mathbf{V}_{t}^{u})$
14:  end for
15:  Compute $z_{t-1}^{s}$ according to Eq. 3.
16:  if $t=1$ then
17:   Compute self-attention maps $\mathcal{A}^{sa}$
18:   $\mathcal{A}^{ca}=\lambda\mathrm{softmax}(\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}^{c}}^{\top}{}}{\sqrt{d}})+(1-\lambda)\mathrm{softmax}(\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}^{u}}^{\top}{}}{\sqrt{d}})$
19:   for $i=I$ to 11 do
20:    $\mathcal{A}^{ca}=\mathcal{A}^{sa}\cdot\mathcal{A}^{ca}$
21:   end for
22:   $\hat{m}=\text{Binarize}(\mathcal{A}^{ca})$
23:  end if
24:end for
25: Use VAE image decoder to recover $z_{0}$ into the image $x^{s}$.
26:return $x^{s}$, $\hat{m}$
@@ IV-B Datasets and Metrics
Industrial inspection. We evaluate Crucial-Diff on the MVTec dataset [56], which contains 3,629 normal training images and a test set of 467 normal and 1,258 anomaly images with mask labels. Additionally, we evaluate our method on the Real-IAD dataset [57]. This dataset contains 150K images covering 30 different object categories, providing more diverse and challenging real-world industrial scenarios. Following [15], one-third of the anomaly images are used for training, while the rest is used to test downstream tasks. Generation performance is evaluated with Intra-cluster pairwise LPIPS distance (IC-LPIPS) [58], Inception Score (IS) [59] and Kernel Inception Distance (KID) [60]. Detection performance is evaluated with Average Precision (AP) and maximum F1 score (F1-MAX). Computational efficiency is measured by metric Time, which means total training hours for all targets.
@@
Figure 4: Visualization comparison of synthetic images generated by Crucial-Diff and other methods. (a) is MVTec result, (b) is Real-IAD dataset result and (c) is polyp dataset result with the annotations of synthetic targets at bottom right of each image. Our generated samples exhibit features like seamless background fusion, surface irregularities or subtle color variations that effectively disrupt downstream models and significantly increase the difficulty of target distinction.
@@ IV-C Qualitative Comparison
IV-C1 Comparison of Synthetic Anomalies
As shown in Fig. 4(a), in MVTec dataset, previous methods, such as AnomalyDiffusion [15], tend to generate targets that are visually distinct and easily recognizable, limiting their effectiveness in improvement of downstream models. In contrast, Crucial-Diff excels at generating subtle anomalies, such as faint cuts and barely visible breaks. These subtle targets are easily overlooked by downstream tasks as training data can effectively improve detection results. This advantage is further validated on the Real-IAD dataset, as illustrated in Fig. 4(b). While other methods, particularly DFMGAN [23], struggle to restore the original appearance of objects, our method successfully generates realistic and subtle damage.
In Fig. 4(c), on the polyp dataset, Crucial-Diff synthesizes polyps with blurry boundaries and low contrast, allowing them to blend naturally into the surrounding tissue. This not only improves visual realism but also enhances the model’s applicability in clinical scenarios. In comparison, previous methods often produce polyps with sharp edges and high contrast. While such synthetic samples are easier to detect, they lack anatomical plausibility and reduce practical utility.
IV-C2 Distribution Analysis of Synthetic Anomalies